OpenStudy (anonymous):

Kevin and Randy Muise have a jar containing 53 coins, all of which are either quarters or nickels. The total value of the coins in the jar is $10.25. How many of each type of coin do they have? The jar contains ____ quarters. The jar contains ____ nickels.

OpenStudy (anonymous):

how much is 1 quarter? how much is 1 nickel?

OpenStudy (anonymous):

this question uses 2 equations and 2 unknowns :D

OpenStudy (texaschic101):

q + n = 53 --> q = 53 - n .25q + .05n = 10.25 now sub 53 - n in for q in the 2nd equation and solve for n. Once n is found, sub that into the 1st equation and solve for q. You can then substitute your know answers into the 2nd equation and if the answer comes out equal, you got it right, however, if the answer does not come out equal, then there was a mistake made.